OpenStudy (anonymous):

Determine whether the geometric series is convergent or divergent. If it is convergent, find its sum. pi^2/(3^n+1)

OpenStudy (anonymous):

Having trouble with these concepts for some reason..

OpenStudy (anonymous):

Im sorry its supposed to be pi^n

OpenStudy (zzr0ck3r):

pi^n/(3^(n+1)) = (1/3) * (pi/3)^n correct?

OpenStudy (anonymous):

im not sure how you got that..

OpenStudy (zzr0ck3r):

(1/3)(1/3^n) = 1/(3^(n+1))

OpenStudy (zzr0ck3r):

a^b*a^c = a^(b+c)

OpenStudy (anonymous):

ok yes.

OpenStudy (zzr0ck3r):

what is the common ratio for (1/3) * (pi/3)^n ?

OpenStudy (zzr0ck3r):

a(b)^n then b is common ratio

OpenStudy (anonymous):

pi/3^n? and a =1/3?

OpenStudy (zzr0ck3r):

ratio is pi/3

OpenStudy (zzr0ck3r):

pi/3 > 1 so it does not converge

OpenStudy (anonymous):

oh ok right.

OpenStudy (zzr0ck3r):

this was all about the algebra to get it into the form a(b)^n get used to factoring things in/out of the denominator to increase/decrease the exponent

OpenStudy (zzr0ck3r):

its a handy "trick"

OpenStudy (anonymous):

oh ok...im sorry...how did you get pi/3^n again?

OpenStudy (amistre64):

is that n+1 an exponent \(k^{(n+1)}\)? or is it \(k^{(n)}+1\)

OpenStudy (anonymous):

its the first one.

OpenStudy (amistre64):

then zzr has this taken care of :)

OpenStudy (zzr0ck3r):

(pi)^n/3^(n+1) = (pi)^n * (1/3^(n+1)) = (pi)^n * (1/3) * (1/3^n) = (1/3) * (pi)^n/3^n = (1/3) (pi/3)^n

OpenStudy (anonymous):

oh ok I got it! I was just having alittle trouble deciphering what you were saying...I get it now!

OpenStudy (zzr0ck3r):

\[\frac{\pi^n}{3^{n+1}}= (1/3)\frac{\pi^n}{3^n}= (1/3)(\frac{\pi}{3})^n\]

OpenStudy (anonymous):

i get it! Thanks alot zzr

OpenStudy (zzr0ck3r):

ok sorry, the equation editor was not working for me but is now:)

OpenStudy (zzr0ck3r):

np